OpenStudy (anonymous):

I'll Medal AND FAN!!! What is 2 and 2 thirds minus 3 fifths??

OpenStudy (anonymous):



OpenStudy (anonymous):

\[2\tfrac{2}{3}=\frac{8}{3}\] is a start

OpenStudy (anonymous):

now compute \[\frac{8}{3}-\frac{3}{5}\]

OpenStudy (anonymous):

8-3 is 5

OpenStudy (anonymous):

@satellite73 Could you walk me through it?

OpenStudy (anonymous):

You need to find a common denominator

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

15 @jtryon

OpenStudy (anonymous):

5/15

OpenStudy (anonymous):